# João Mattos, Debolina Halder Lina, Arlei Silva

### v1

@@ 5.1 Hidden Bias in Dyadic Fairness
A) UGE B) EDITS C) FairAdj D) GRAPHAIR E) FairWalk F) DELTR G) DetConstSort H) MORAL
Figure 3: Comparison of the proportions of each pair type in the top-$100$ predictions from each method in the six real-world datasets considered. We exhibit two references: the original graph distribution and the optimal dyadic fairness reference. $E_{s\text{-}s}$ pairs are colored in blue, while $E_{s^{\prime}\text{-}s}$ and $E_{s^{\prime}\text{-}s^{\prime}}$ are colored in black and red, respectively. The purple color in the dyadic fairness reference bar corresponds to the sum of the proportions of $E_{s^{\prime}\text{-}s^{\prime}}$ and $E_{s\text{-}s}$ pairs. MORAL obtains the top-$100$ that better represents the original graph distribution, while all baselines present significant deviations from the target distribution. Bars are absent in case of an out-of-memory error.
We first aim to answer RQ1, demonstrating the limitations of dyadic fairness for link prediction. We analyze the distributions of types of pairs in the top-$k$ of each method, independent from the order of the elements. In Figure 3, we compare the proportions of pairs in the top-$k$ ($k=100$) of each baseline. Our approach obtains the closest approximation to the target distribution, while other methods overestimate one pair type at the detriment of another, despite obtaining low values of $\Delta_{DP}$ . This result demonstrates the limitation of $\Delta_{DP}$ in detecting underrepresented aggregated groups, and sheds light on the importance of adhering to Property 1.
5.2 Demographic Parity Gap
To answer RQ2, we demonstrate the effect of ranking on fairness in link prediction by fixing the proportion of each pair type. First, we compute the required proportions of each pair type for optimal $\Delta_{DP}$ . Then, we fix these proportions, but consider the worst and best possible permutations of these candidate pairs in terms of NDKL. To isolate the fairness component from the utility measurement, we assume in this experiment that the pairs being ranked are all positive, meaning that regardless of the permutation of the pairs, both rankings will obtain $prec@k=100\%$. We expose in Appendix B the NDKL gap for varying-sized rankings.
Despite relevant exposure bias between the worst and best rankings, the $\Delta_{DP}$ measurement is the same. This indicates the necessity of incorporating Property 2 into fair link prediction metrics.
5.3 Baselines Ranking Comparison
@@
We compare our method against UGE, EDITS, FairAdj, GRAPHAIR, FairWalk, DELTR and DetConstSort. We adopt NDKL as fairness metric and prec@1000 as utility metric. For EDITS, we use the dot product between the node embeddings obtained by training a GCN on the fair graph generated as the pair scores. For GRAPHAIR, UGE and FairWalk, the scores are obtained through the dot product of the fair node embeddings outputted.
We show our results on Table 2 and Appendix G. MORAL achieves the fairest rankings while still maintaining high link prediction performance. Both periphery and community graph types are challenging for all methods. We highlight the large fairness improvements obtained by our approach across all datasets, particularly in Credit, German and NBA.
@@ 6 Related Work
We situate MORAL within two major areas: fair graph representation learning and fair ranking. Our approach is an example of a post-processing method, which have proven effective in fair ranking tasks (xian2023fair; tifrea2023frappe; gorantla_problem_2021; zehlike_fair_2017; li2021user), and aligns with decoupling classifiers for group fairness, which has been successfully explored in the past (dwork_decoupled_2018).
@@ 6.1 Fair Graph Representation Learning
In-processing. FairGNN (dai_say_2021), NIFTY (agarwal_towards_2021), GRAPHAIR (ling_learning_2023), FairVGNN (wang_improving_2022) and UGE (wang2022unbiased) integrate fairness during training using adversarial learning or graph augmentations. These methods assume fairness can be enforced at the node embedding level, which may encounter expressivity power limitations, and not generalize to pairwise tasks like link prediction.
@@ 6.2 Fair Ranking and Information Retrieval
Post-processing approaches like FA*IR (zehlike_fair_2017) and DetConstSort (geyik_fairness-aware_2019) re-rank outputs to satisfy fairness constraints. Though model-agnostic, they can be computationally intensive and inflexible. In-processing ranking methods, such as DELTR (zehlike_reducing_2020), policy learning (singh_policy_2019; yadav_policy-gradient_2021), and constrained optimization via SPOFR (kotary_end--end_2021), jointly optimize fairness and relevance. However, they typically target dyadic ranking tasks in learn-to-rank settings, which is misaligned with the structural nature of link prediction task considered in this work.
@@ 7 Conclusion
Inspired by the Information Retrieval literature, we formulate two properties to detect bias issues that can be hidden when only considering demographic parity. Our experiments demonstrate the effect of the absence of each of these properties. Finally we propose a simple score post-processing method that mitigates bias in all datasets considered. We hope our efforts can push the boundaries of fairness research in tasks such as link prediction and can be a stepping stone for future advancements in the field.
@@ Appendix G Additional Performance Results
In addition to the $prec@1000$ results exposed in Subsection 5.3, we also provide performance results on $prec@100$ in Table 3.

### v2

@@ 5.1 Hidden Bias in Dyadic Fairness
A) UGE B) EDITS C) FairAdj D) FairEGM E) FairLP F) GRAPHAIR G) FairWalk H) DELTR I) DetConstSort J) MORAL
Figure 3: Proportions of pair types in the top-$100$ predictions by method, compared against the original graph distribution and an optimal dyadic fairness reference. Colors: $E_{s\text{-}s}$ (blue), $E_{s^{\prime}\text{-}s}$ (black), and $E_{s^{\prime}\text{-}s^{\prime}}$ (red). In the dyadic fairness reference, purple represents the combined proportion of $E_{s^{\prime}\text{-}s^{\prime}}$ and $E_{s\text{-}s}$ pairs. Missing bars indicate an OOM error.
@@
FairEGM NDKL 0.09 $\pm$ 0.01 0.11 $\pm$ 0.00 0.05 $\pm$ 0.01 0.07 $\pm$ 0.01 OOM OOM
$prec@1000$ 0.97 $\pm$ 0.00 1.00 $\pm$ 0.00 0.62 $\pm$ 0.00 0.60 $\pm$ 0.01 OOM OOM
FairLP NDKL 0.18 $\pm$ 0.00 OOM 0.06 $\pm$ 0.00 0.20 $\pm$ 0.00 OOM OOM
$prec@1000$ 0.99 $\pm$ 0.00 OOM 0.97 $\pm$ 0.00 0.86 $\pm$ 0.00 OOM OOM
@@
We first aim to answer RQ1, demonstrating the limitations of dyadic fairness for link prediction. We analyze the distributions of types of pairs in the top-$k$ of each method, independent of the order of the elements. In Figure 3, we compare the proportions of pairs in the top-$k$ ($k=100$) of each baseline. Our approach obtains the closest approximation to the target distribution, while other methods overestimate one pair type at the detriment of another, despite obtaining low values of $\Delta_{DP}$ . This result demonstrates the limitation of $\Delta_{DP}$ in detecting underrepresented aggregated groups, and sheds light on the importance of adhering to Property 1.
5.2 Demographic Parity Gap
To answer RQ2, we demonstrate the effect of ranking on fairness in link prediction by fixing the proportion of each pair type. First, we compute the required proportions of each pair type for optimal $\Delta_{DP}$ . Then, we fix these proportions, but consider the worst and best possible permutations of these candidate pairs in terms of NDKL. To isolate the fairness component from the utility measurement, we assume in this experiment that the pairs being ranked are all positive, meaning that regardless of the permutation of the pairs, both rankings will obtain $prec@k=100\%$. We expose the NDKL gap for varying-sized rankings in the Appendix B. Despite relevant exposure bias between the worst and best rankings, the value of $\Delta_{DP}$ is the same. This indicates the necessity of incorporating Property 2 into fair link prediction metrics.
5.3 Baselines Ranking Comparison
We compare our method against UGE, EDITS, FairAdj, GRAPHAIR, FairWalk, DELTR, and DetConstSort. We adopt NDKL as the fairness metric and prec@1000 as the utility metric. For EDITS, we use the dot product between the node embeddings obtained by training a GCN on the fair graph generated as the pair scores. For GRAPHAIR, UGE, and FairWalk, the scores are obtained through the dot product of the fair node embeddings outputted.
We show our results in Table 2. MORAL achieves the fairest rankings while still maintaining high link prediction performance. Both periphery and community graph types are challenging for all methods. We highlight the large fairness improvements obtained by our approach across all datasets, particularly in Credit, German, and NBA.
@@ 6 Related Work
We situate MORAL within two major areas: fair graph representation learning and fair ranking. Our approach is an example of a post-processing method, which has proven effective in fair ranking tasks (xian2023fair; tifrea2023frappe; gorantla_problem_2021; zehlike_fair_2017; li2021user), and aligns with decoupling classifiers for group fairness, which has been successfully explored in the past (dwork_decoupled_2018).
@@ 6.1 Fair Graph Representation Learning
In-processing. FairGNN (dai_say_2021), NIFTY (agarwal_towards_2021), GRAPHAIR (ling_learning_2023), FairVGNN (wang_improving_2022) , and UGE (wang2022unbiased) integrate fairness during training using adversarial learning or graph augmentations. These methods assume fairness can be enforced at the node embedding level, which may encounter expressivity power limitations, and not generalize to pairwise tasks like link prediction.
@@ 6.2 Fair Ranking and Information Retrieval
Post-processing approaches like FA*IR (zehlike_fair_2017) and DetConstSort (geyik_fairness-aware_2019) re-rank outputs to satisfy fairness constraints. Though model-agnostic, they can be computationally intensive and inflexible. In-processing ranking methods, such as DELTR (zehlike_reducing_2020), policy learning (singh_policy_2019; yadav_policy-gradient_2021), and constrained optimization via SPOFR (kotary_end--end_2021), jointly optimize fairness and relevance. However, they typically target dyadic ranking tasks in learn-to-rank settings, which is misaligned with the structural nature of the link prediction task considered in this work.
@@ Appendix G Additional Performance Results
[…]
Method Cora
NIFTY GAE NDKL 0.0540 $\pm$ 0.0030
$prec@1000$ 0.7300 $\pm$ 0.2970
FairWalk NDKL 0.0522 $\pm$ 0.0002
$prec@1000$ 0.5110 $\pm$ 0.0030
DELTR GAE NDKL 0.0484 $\pm$ 0.0001
$prec@1000$ 0.5225 $\pm$ 0.0007
MORAL NDKL 0.0471 $\pm$ 0.0000
$prec@1000$ 0.7229 $\pm$ 0.0010
In addition to the $prec@1000$ and NDKL results exposed in Subsection 5.3, we conduct the evaluation on other performance ($prec@100$, NDCGNDCG, APAP) and fairness ($\Delta_{DP}$) exposed in tables 4, 5, 6, 7 and 8, respectively. We highlight that even though GraphAIR does not obtain the lowest $NDKL$ values, it is the overall most fair outputs considering $\Delta_{DP}$, demonstrating the limitation of the metric.
We simulate a multi-sensitive attribute scenario by considering a multi-label network, i.e. Cora, and label as a sensitive attribute value. In order to avoid the combinatorial cost of training $k\choose 2$ classifiers, we train a single model per sensitive attribute value in a one vs. all classification fashion. We expose the results on a simulated multi-sensitive attribute scenario conducted on Cora on Table 3.
@@
Table 5: Performance comparison (NDKL↓ and AP↑). Best AP per dataset is underlined. Best NDKL per dataset is bold. OOM = method ran out-of-memory.
Method Facebook Credit German NBA Pokec-n Pokec-z
AP 0.93 $\pm$ 0.00 0.85 $\pm$ 0.01 0.60 $\pm$ 0.01 0.56 $\pm$ 0.01 0.63 $\pm$ 0.03 0.63 $\pm$ 0.02
EDITS GAE NDKL 0.21 $\pm$ 0.07 0.04 $\pm$ 0.02 0.08 $\pm$ 0.02 0.08 $\pm$ 0.01 OOM OOM
AP 0.78 $\pm$ 0.01 0.63 $\pm$ 0.01 0.50 $\pm$ 0.00 0.50 $\pm$ 0.00 OOM OOM
AP 0.90 $\pm$ 0.00 0.90 $\pm$ 0.02 0.62 $\pm$ 0.00 0.65 $\pm$ 0.01 0.75 $\pm$ 0.05 0.66 $\pm$ 0.23
FairEGM NDKL 0.09 $\pm$ 0.01 0.11 $\pm$ 0.00 0.05 $\pm$ 0.01 0.07 $\pm$ 0.01 OOM OOM
AP 0.94 $\pm$ 0.00 0.96 $\pm$ 0.00 0.56 $\pm$ 0.00 0.58 $\pm$ 0.01 OOM OOM
FairLP NDKL 0.18 $\pm$ 0.00 OOM 0.06 $\pm$ 0.00 0.20 $\pm$ 0.00 OOM OOM
AP 0.95 $\pm$ 0.00 OOM 0.84 $\pm$ 0.00 0.78 $\pm$ 0.00 OOM OOM
AP 0.91 $\pm$ 0.00 0.98 $\pm$ 0.00 0.86 $\pm$ 0.00 0.54 $\pm$ 0.01 0.86 $\pm$ 0.00 0.86 $\pm$ 0.00
FairAdj NDKL 0.10 $\pm$ 0.05 OOM 0.10 $\pm$ 0.01 0.11 $\pm$ 0.05 OOM OOM
AP 0.55 $\pm$ 0.01 OOM 0.51 $\pm$ 0.00 0.53 $\pm$ 0.01 OOM OOM
DetConstSort GAE NDKL 0.15 $\pm$ 0.00 0.06 $\pm$ 0.00 0.04 $\pm$ 0.00 0.09 $\pm$ 0.00 0.07 $\pm$ 0.00 0.23 $\pm$ 0.00
AP 0.05 $\pm$ 0.00 0.00 $\pm$ 0.00 0.56 $\pm$ 0.00 0.28 $\pm$ 0.00 0.07 $\pm$ 0.00 0.02 $\pm$ 0.00
DELTR GAE NDKL 0.10 $\pm$ 0.03 0.03 $\pm$ 0.00 0.09 $\pm$ 0.06 0.09 $\pm$ 0.02 0.23 $\pm$ 0.23 0.22 $\pm$ 0.20
AP 0.78 $\pm$ 0.04 0.60 $\pm$ 0.02 0.50 $\pm$ 0.03 0.47 $\pm$ 0.14 0.44 $\pm$ 0.08 0.41 $\pm$ 0.06
ThreeClass GAE NDKL 0.04 $\pm$ 0.00 0.01 $\pm$ 0.00 0.03 $\pm$ 0.00 0.03 $\pm$ 0.00 0.02 $\pm$ 0.00 0.03 $\pm$ 0.00
AP 0.94 $\pm$ 0.00 1.00 $\pm$ 0.00 0.96 $\pm$ 0.01 0.80 $\pm$ 0.00 0.98 $\pm$ 0.00 0.98 $\pm$ 0.00
Table 6: Performance comparison (hits@1000hits@1000↑ and NDKL↓). Best $hits@1000$ per dataset is underlined. Best NDKL per dataset is bold. OOM = method ran out-of-memory.
Method Facebook Credit German NBA Pokec-n Pokec-z
$hits@1000$ 0.92 $\pm$ 0.00 0.51 $\pm$ 0.01 0.36 $\pm$ 0.01 0.58 $\pm$ 0.03 0.05 $\pm$ 0.01 0.04 $\pm$ 0.01
EDITS GAE NDKL 0.21 $\pm$ 0.07 0.04 $\pm$ 0.02 0.08 $\pm$ 0.02 0.08 $\pm$ 0.01 OOM OOM
$hits@1000$ 0.79 $\pm$ 0.21 0.56 $\pm$ 0.38 0.48 $\pm$ 0.29 0.36 $\pm$ 0.15 OOM OOM
$hits@1000$ 0.85 $\pm$ 0.01 0.49 $\pm$ 0.05 0.39 $\pm$ 0.01 0.69 $\pm$ 0.02 0.11 $\pm$ 0.05 0.08 $\pm$ 0.12
FairEGM NDKL 0.09 $\pm$ 0.01 0.11 $\pm$ 0.00 0.05 $\pm$ 0.01 0.07 $\pm$ 0.01 OOM OOM
$hits@1000$ 0.95 $\pm$ 0.00 0.76 $\pm$ 0.01 0.29 $\pm$ 0.00 0.62 $\pm$ 0.02 OOM OOM
FairLP NDKL 0.18 $\pm$ 0.00 OOM 0.06 $\pm$ 0.00 0.20 $\pm$ 0.00 OOM OOM
$hits@1000$ 0.94 $\pm$ 0.00 OOM 0.77 $\pm$ 0.00 0.83 $\pm$ 0.00 OOM OOM
$hits@1000$ 0.89 $\pm$ 0.01 0.92 $\pm$ 0.00 0.80 $\pm$ 0.00 0.54 $\pm$ 0.02 0.33 $\pm$ 0.00 0.32 $\pm$ 0.00
FairAdj NDKL 0.10 $\pm$ 0.05 OOM 0.10 $\pm$ 0.01 0.11 $\pm$ 0.05 OOM OOM
$hits@1000$ 0.16 $\pm$ 0.01 OOM 0.27 $\pm$ 0.01 0.51 $\pm$ 0.03 OOM OOM
DetConstSort GAE NDKL 0.15 $\pm$ 0.00 0.06 $\pm$ 0.00 0.04 $\pm$ 0.00 0.09 $\pm$ 0.00 0.07 $\pm$ 0.00 0.23 $\pm$ 0.00
$hits@1000$ 1.00 $\pm$ 0.00 0.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00
DELTR GAE NDKL 0.10 $\pm$ 0.03 0.03 $\pm$ 0.00 0.09 $\pm$ 0.06 0.09 $\pm$ 0.02 0.23 $\pm$ 0.23 0.22 $\pm$ 0.20
$hits@1000$ 0.62 $\pm$ 0.05 0.00 $\pm$ 0.00 0.12 $\pm$ 0.17 0.39 $\pm$ 0.26 0.01 $\pm$ 0.00 0.01 $\pm$ 0.01
ThreeClass GAE NDKL 0.04 $\pm$ 0.00 0.01 $\pm$ 0.00 0.03 $\pm$ 0.00 0.03 $\pm$ 0.00 0.02 $\pm$ 0.00 0.03 $\pm$ 0.00
$hits@1000$ 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00
Table 7: Performance comparison (NDCG↑ and NDKL↓). Best NDCG per dataset is underlined. Best NDKL per dataset is bold. OOM = method ran out-of-memory.
Method Facebook Credit German NBA Pokec-n Pokec-z
NDCG 0.99 $\pm$ 0.00 0.98 $\pm$ 0.00 0.94 $\pm$ 0.00 0.92 $\pm$ 0.00 0.96 $\pm$ 0.00 0.96 $\pm$ 0.00
EDITS GAE NDKL 0.21 $\pm$ 0.07 0.04 $\pm$ 0.02 0.08 $\pm$ 0.02 0.08 $\pm$ 0.01 OOM OOM
NDCG 0.97 $\pm$ 0.00 0.95 $\pm$ 0.00 0.91 $\pm$ 0.00 0.90 $\pm$ 0.00 OOM OOM
NDCG 0.99 $\pm$ 0.00 0.99 $\pm$ 0.00 0.94 $\pm$ 0.00 0.94 $\pm$ 0.00 0.97 $\pm$ 0.01 0.96 $\pm$ 0.03
FairEGM NDKL 0.09 $\pm$ 0.01 0.11 $\pm$ 0.00 0.05 $\pm$ 0.01 0.07 $\pm$ 0.01 OOM OOM
NDCG 0.99 $\pm$ 0.00 1.00 $\pm$ 0.00 0.93 $\pm$ 0.00 0.92 $\pm$ 0.00 OOM OOM
FairLP NDKL 0.18 $\pm$ 0.00 OOM 0.06 $\pm$ 0.00 0.20 $\pm$ 0.00 OOM OOM
NDCG 0.99 $\pm$ 0.00 OOM 0.98 $\pm$ 0.00 0.97 $\pm$ 0.00 OOM OOM
NDCG 0.99 $\pm$ 0.00 1.00 $\pm$ 0.00 0.98 $\pm$ 0.00 0.91 $\pm$ 0.00 0.99 $\pm$ 0.00 0.99 $\pm$ 0.00
FairAdj NDKL 0.10 $\pm$ 0.05 OOM 0.10 $\pm$ 0.01 0.11 $\pm$ 0.05 OOM OOM
NDCG 0.91 $\pm$ 0.00 OOM 0.91 $\pm$ 0.00 0.91 $\pm$ 0.00 OOM OOM
DetConstSort GAE NDKL 0.15 $\pm$ 0.00 0.06 $\pm$ 0.00 0.04 $\pm$ 0.00 0.09 $\pm$ 0.00 0.07 $\pm$ 0.00 0.23 $\pm$ 0.00
NDCG 0.23 $\pm$ 0.00 0.00 $\pm$ 0.00 0.89 $\pm$ 0.00 0.77 $\pm$ 0.00 0.51 $\pm$ 0.00 0.28 $\pm$ 0.00
DELTR GAE NDKL 0.10 $\pm$ 0.03 0.03 $\pm$ 0.00 0.09 $\pm$ 0.06 0.09 $\pm$ 0.02 0.23 $\pm$ 0.23 0.22 $\pm$ 0.20
NDCG 0.97 $\pm$ 0.01 0.94 $\pm$ 0.00 0.91 $\pm$ 0.01 0.88 $\pm$ 0.06 0.93 $\pm$ 0.02 0.92 $\pm$ 0.02
ThreeClass GAE NDKL 0.04 $\pm$ 0.00 0.01 $\pm$ 0.00 0.03 $\pm$ 0.00 0.03 $\pm$ 0.00 0.02 $\pm$ 0.00 0.03 $\pm$ 0.00
NDCG 0.99 $\pm$ 0.00 1.00 $\pm$ 0.00 0.99 $\pm$ 0.00 0.97 $\pm$ 0.00 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00
Table 8: Fairness comparison using $\Delta_{DP}$↓ and Precision↑. Best (lowest) $\Delta_{DP}$ per dataset is bold. Best (highest) Precision per dataset is underlined. OOM = method ran out-of-memory.
Method Facebook Credit German NBA Pokec-n Pokec-z
UGE $\Delta_{DP}$ 0.00 $\pm$ 0.00 0.03 $\pm$ 0.00 0.01 $\pm$ 0.00 0.00 $\pm$ 0.00 0.02 $\pm$ 0.00 0.02 $\pm$ 0.00
Precision 0.97 $\pm$ 0.00 1.00 $\pm$ 0.00 0.69 $\pm$ 0.01 0.58 $\pm$ 0.01 0.90 $\pm$ 0.04 0.91 $\pm$ 0.04
EDITS GAE $\Delta_{DP}$ 0.01 $\pm$ 0.00 0.06 $\pm$ 0.04 0.01 $\pm$ 0.01 0.00 $\pm$ 0.00 OOM OOM
Precision 0.96 $\pm$ 0.00 0.36 $\pm$ 0.17 0.42 $\pm$ 0.20 0.49 $\pm$ 0.00 OOM OOM
GRAPHAIR $\Delta_{DP}$ 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.01 $\pm$ 0.01
Precision 0.96 $\pm$ 0.01 1.00 $\pm$ 0.00 0.73 $\pm$ 0.01 0.69 $\pm$ 0.01 0.97 $\pm$ 0.03 1.00 $\pm$ 0.00
FairEGM $\Delta_{DP}$ 0.01 $\pm$ 0.00 0.02 $\pm$ 0.00 0.01 $\pm$ 0.00 0.03 $\pm$ 0.00 OOM OOM
Precision 0.97 $\pm$ 0.00 1.00 $\pm$ 0.00 0.62 $\pm$ 0.00 0.60 $\pm$ 0.01 OOM OOM
FairLP $\Delta_{DP}$ 0.00 $\pm$ 0.00 OOM 0.01 $\pm$ 0.00 0.01 $\pm$ 0.00 OOM OOM
Precision 0.99 $\pm$ 0.00 OOM 0.97 $\pm$ 0.00 0.86 $\pm$ 0.00 OOM OOM
FairWalk $\Delta_{DP}$ 0.00 $\pm$ 0.00 0.01 $\pm$ 0.00 0.02 $\pm$ 0.00 0.00 $\pm$ 0.00 0.03 $\pm$ 0.00 0.03 $\pm$ 0.00
Precision 0.96 $\pm$ 0.00 1.00 $\pm$ 0.00 0.94 $\pm$ 0.00 0.55 $\pm$ 0.01 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00
FairAdj $\Delta_{DP}$ 0.15 $\pm$ 0.12 OOM 0.52 $\pm$ 0.24 0.30 $\pm$ 0.23 OOM OOM
Precision 0.42 $\pm$ 0.01 OOM 0.54 $\pm$ 0.01 0.50 $\pm$ 0.01 OOM OOM
DetConstSort GAE $\Delta_{DP}$ 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.01 $\pm$ 0.00 0.01 $\pm$ 0.00 0.02 $\pm$ 0.00
Precision 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.55 $\pm$ 0.00 0.21 $\pm$ 0.00 0.07 $\pm$ 0.00 0.01 $\pm$ 0.00
DELTR GAE $\Delta_{DP}$ 0.00 $\pm$ 0.00 0.03 $\pm$ 0.02 0.03 $\pm$ 0.01 0.01 $\pm$ 0.02 0.05 $\pm$ 0.05 0.04 $\pm$ 0.03
Precision 0.91 $\pm$ 0.05 0.56 $\pm$ 0.29 0.31 $\pm$ 0.44 0.43 $\pm$ 0.24 0.65 $\pm$ 0.01 0.48 $\pm$ 0.28
ThreeClass GAE $\Delta_{DP}$ 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.01 $\pm$ 0.00 0.07 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00
Precision 0.95 $\pm$ 0.01 1.00 $\pm$ 0.00 0.96 $\pm$ 0.00 0.80 $\pm$ 0.00 0.98 $\pm$ 0.00 0.98 $\pm$ 0.00